# Completeness of coalition logics with seriality, independence of agents, or determinism

### v1

@@ 3.3.1 Blueprints and their realization
    Then, for some $\phi\in\mathtt{list}_{\mathrm{C}}({\sigma_{\mathrm{C}}})$, $\phi\vDash_{\mathsf{MCL}+\mathsf{X}}\psi$. Then for some ${\sigma_{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}$, ${\sigma_{\mathrm{C}}}\subseteq{\sigma_{\mathtt{AG}}}$ and $\phi\in\mathtt{list}_{\mathtt{AG}}({\sigma_{\mathtt{AG}}})$. Then ${\sigma_{\mathtt{AG}}}\in\mathtt{PJA}_{\mathtt{AG}}$. Then the pointed $\mathsf{X}$-model $(\mathtt{M}^{\sigma_{\mathtt{AG}}}_{\phi},s^{\sigma_{\mathtt{AG}}}_{\phi})$ is defined and $\mathtt{M}^{\sigma_{\mathtt{AG}}}_{\phi},s^{\sigma_{\mathtt{AG}}}_{\phi}\Vdash\phi$. Then $\mathtt{M},s^{\sigma_{\mathtt{AG}}}_{\phi}\Vdash\phi$. Then, $\mathtt{M},s^{\sigma_{\mathtt{AG}}}_{\phi}\Vdash\psi$. Note $s^{\sigma_{\mathtt{AG}}}_{\phi}\in\mathsf{out}_{\mathtt{AG}}(s_{0},\sigma_{\mathtt{AG}})$. Then $s^{\sigma_{\mathtt{AG}}}_{\phi}\in\mathsf{out}_{\mathrm{C}}(s_{0},\sigma_{\mathrm{C}})$. Therefore, ${\sigma_{\mathrm{C}}}\ltimes_{(\mathtt{M},s_{0})}\psi$.
3.3.2 The lemma
Let $\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\to\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$ be a standard disjunction. We use $\mathbb{J}_{0}$ to indicate $\{j\in\mathbb{J}\mid\mathrm{B}_{j}=\mathtt{AG}\}$.
@@ 3.3.2 The lemma
Let $\mathsf{X}\in\mathrm{ES}$ and $\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\to\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$ be a standard disjunction. Then, if $\models_{{\mathsf{MCL}+\mathsf{X}}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\to\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$, then the following condition, called the $\mathsf{X}$-validity-reduction-condition, is met:
@@
Assume the $\mathsf{X}$-validity-reduction-condition is not met. Then, it can be verified that the following condition, called the $\mathsf{X}$-satisfiability-reduction-condition, is met:
@@
Let $\gamma^{\prime}$ be an $\mathsf{X}$-satisfiable elementary conjunction, equivalent to $\neg\gamma$. It suffices to show the standard conjunction $\gamma^{\prime}\wedge\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\wedge\bigwedge_{j\in\mathbb{J}}\neg\langle[\mathrm{B}_{j}]\rangle\psi_{j}$ is $\mathsf{X}$-satisfiable.
In the sequel, first, we define a blueprint, from which we can get a pointed $\mathsf{X}$-model; second, we show the pointed $\mathsf{X}$-model satisfies $\gamma^{\prime}\wedge\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\wedge\bigwedge_{j\in\mathbb{J}}\neg\langle[\mathrm{B}_{j}]\rangle\psi_{j}$.
@@
Define a blueprint $\mathtt{BP}=(\mathtt{AG},\mathtt{AC}_{0},\mathtt{list}_{\mathtt{AG}})$ as follows:
@@
    $\begin{cases}\Big{\{}\bigwedge_{i\in\mathtt{support}({\sigma_{\mathtt{AG}}})}\phi_{i}\wedge\neg\psi_{j}\mid j\in\mathbb{J}\text{ and }\bigcup_{i\in\mathtt{support}({\sigma_{\mathtt{AG}}})}\mathrm{A}_{i}\subseteq\mathrm{B}_{j}\Big{\}}&\parbox[t]{120.00018pt}{ \text{if $\mathtt{support}({\sigma_{\mathtt{AG}}})$ is $\mathsf{X}$-neat} }\\ \emptyset&\text{otherwise}\end{cases}$
@@
    $\begin{cases}\Big{\{}\bigwedge_{i\in\mathtt{support}({\sigma_{\mathtt{AG}}})}\phi_{i}\wedge\neg\psi_{\mathtt{impeach}({\sigma_{\mathtt{AG}}})}\wedge\bigwedge_{k\in\mathbb{J}_{0}}\neg\psi_{k}\Big{\}}&\parbox[t]{200.0003pt}{ \text{if $\mathtt{support}({\sigma_{\mathtt{AG}}})$ is $\mathsf{X}$-neat and} }\\ &\parbox[t]{200.0003pt}{ \text{$\bigcup_{i\in\mathtt{support}({\sigma_{\mathtt{AG}}})}\mathrm{A}_{i}\subseteq\mathrm{B}_{\mathtt{impeach}({\sigma_{\mathtt{AG}}})}$} }\vspace{15pt}\\ \Big{\{}\bigwedge_{i\in\mathtt{support}({\sigma_{\mathtt{AG}}})}\phi_{i}\wedge\neg\psi_{0}\wedge\bigwedge_{k\in\mathbb{J}_{0}}\neg\psi_{k}\Big{\}}&\parbox[t]{200.0003pt}{ \text{if $\mathtt{support}({\sigma_{\mathtt{AG}}})$ is $\mathsf{X}$-neat and} }\\ &\parbox[t]{200.0003pt}{ \text{$\bigcup_{i\in\mathtt{support}({\sigma_{\mathtt{AG}}})}\mathrm{A}_{i}\not\subseteq\mathrm{B}_{\mathtt{impeach}({\sigma_{\mathtt{AG}}})}$} }\vspace{15pt}\\ \emptyset&\text{otherwise}\end{cases}$
@@
    • Assume $\mathsf{D}\notin\mathsf{X}$. Then there is $j\in\mathbb{J}$ such that $\chi=\bigwedge_{i\in\mathtt{support}({\sigma_{\mathtt{AG}}})}\phi_{i}\wedge\neg\psi_{j}$, where $\mathtt{support}({\sigma_{\mathtt{AG}}})$ is $\mathsf{X}$-neat and $\bigcup_{i\in\mathtt{support}({\sigma_{\mathtt{AG}}})}\mathrm{A}_{i}\subseteq\mathrm{B}_{j}$. By the $\mathsf{X}$-satisfiability-reduction-condition, $\chi$ is $\mathsf{X}$-satisfiable.
@@
      • Assume $\mathtt{support}({\sigma_{\mathtt{AG}}})=\emptyset$. We want to show this case is impossible. By Claim (3), $\mathbb{I}=\emptyset$. Therefore, $\mathtt{support}(\lambda_{\mathtt{AG}})=\emptyset$ for all $\lambda_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$. Note ${\sigma_{\mathrm{C}}}\in\mathtt{PJA}_{\mathrm{C}}$. Then $\mathtt{list}_{\mathrm{C}}(\sigma_{\mathrm{C}})\neq\emptyset$. Then there is $\lambda_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\lambda_{\mathtt{AG}}$ and $\mathtt{list}_{\mathtt{AG}}(\lambda_{\mathtt{AG}})\neq\emptyset$. By Claim (4), $\mathtt{support}(\lambda_{\mathtt{AG}})$ is $\mathsf{X}$-neat. Note $\mathsf{S}\notin\mathsf{X}$. Then $\mathtt{support}(\lambda_{\mathtt{AG}})\neq\emptyset$. We have a contradiction.
@@
By Theorem 1, there is a pointed $\mathsf{X}$-model $(\mathtt{M},s)$, where $\mathtt{M}=(\mathtt{ST},\mathtt{AG},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathtt{AP},\mathsf{label})$, realizing $\mathtt{BP}$ and $\gamma^{\prime}$. We claim $\mathtt{M},s\Vdash\gamma^{\prime}\land\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\wedge\bigwedge_{j\in\mathbb{J}}\neg\langle[\mathrm{B}_{j}]\rangle\psi_{j}$.
Let $i\in\mathbb{I}$. We want to show $\mathtt{M},s\Vdash\langle[\mathrm{A}_{i}]\rangle\phi_{i}$. Let ${\sigma_{\mathrm{A}_{i}}}\in\mathtt{JA}_{\mathrm{A}_{i}}$ such that ${{\sigma_{\mathrm{A}_{i}}}(a)}=i$ for all $a\in\mathrm{A}_{i}$. It suffices to show $\sigma_{\mathrm{A}_{i}}\in\mathtt{PJA}_{\mathrm{A}_{i}}$ and ${\sigma_{\mathrm{A}_{i}}}\vartriangleright_{(\mathtt{M},s)}\phi_{i}$.
@@
Let $j\in\mathbb{J}$. We want to show $\mathtt{M},s\Vdash\neg\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Let ${\sigma_{\mathrm{B}_{j}}}\in\mathtt{PJA}_{\mathrm{B}_{j}}$. It suffices to show ${\sigma_{\mathrm{B}_{j}}}\ltimes_{(\mathtt{M},s)}\neg\psi_{j}$.
Then, $\mathtt{list}_{\mathrm{B}_{i}}({\sigma_{\mathrm{B}_{j}}})\neq\emptyset$. Then there is $\lambda_{\mathtt{AG}}\in\mathtt{PJA}_{\mathtt{AG}}$ such that ${\sigma_{\mathrm{B}_{j}}}\subseteq\lambda_{\mathtt{AG}}$ and $\mathtt{list}_{\mathtt{AG}}(\lambda_{\mathtt{AG}})\neq\emptyset$. By Claim (4), $\mathtt{support}(\lambda_{\mathtt{AG}})$ is $\mathsf{X}$-neat. By Claim (2) and Fact 2, $\mathtt{support}({\sigma_{\mathrm{B}_{j}}})$ is $\mathsf{X}$-neat.
@@
    • Assume $\mathrm{A}_{i}\nsubseteq\mathrm{B}_{j}$. We want to show this case is impossible. Let $\overline{b}\in\mathrm{A}_{i}-\mathrm{B}_{j}$. Then $\overline{b}\in\mathtt{AG}-\mathrm{B}_{j}$. Then ${{\sigma_{\mathtt{AG}}}(\overline{b})}=j$. Note $\overline{b}\in\mathrm{A}_{i}$. Then ${{\sigma_{\mathtt{AG}}}(\overline{b})}=i$. Then $j=i$. Note $i\in\mathbb{I}$ and $j\in\mathbb{J}$. Then $\mathbb{I}\cap\mathbb{J}\neq\emptyset$. We have a contradiction.

### v2

@@ 3.4 Blueprints and their realization
    Then, for some $\phi\in\mathtt{list}_{\mathrm{C}}({\sigma_{\mathrm{C}}})$, $\phi\vDash_{\mathsf{MCL}+\mathsf{X}}\psi$. This implies that for some ${\sigma_{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}$, ${\sigma_{\mathrm{C}}}\subseteq{\sigma_{\mathtt{AG}}}$ and $\phi\in\mathtt{list}_{\mathtt{AG}}({\sigma_{\mathtt{AG}}})$. Thus, ${\sigma_{\mathtt{AG}}}\in\mathtt{PJA}_{\mathtt{AG}}$. Then the pointed $\mathsf{X}$-model $(\mathtt{M}^{\sigma_{\mathtt{AG}}}_{\phi},s^{\sigma_{\mathtt{AG}}}_{\phi})$ is defined and $\mathtt{M}^{\sigma_{\mathtt{AG}}}_{\phi},s^{\sigma_{\mathtt{AG}}}_{\phi}\Vdash\phi$. Then $\mathtt{M},s^{\sigma_{\mathtt{AG}}}_{\phi}\Vdash\phi$. Then, $\mathtt{M},s^{\sigma_{\mathtt{AG}}}_{\phi}\Vdash\psi$. Note $s^{\sigma_{\mathtt{AG}}}_{\phi}\in\mathsf{out}_{\mathtt{AG}}(s_{0},\sigma_{\mathtt{AG}})$. Then $s^{\sigma_{\mathtt{AG}}}_{\phi}\in\mathsf{out}_{\mathrm{C}}(s_{0},\sigma_{\mathrm{C}})$. Therefore, ${\sigma_{\mathrm{C}}}\ltimes_{(\mathtt{M},s_{0})}\psi$.
3.5 Downward validity lemma
The following notion will be used in stating the downward validity lemma, and also the upward derivability lemma in the next subsection.
Definition 16 (Neat sets of negative indices).
Let $\mathsf{X}\in\mathrm{ES}$ and $\gamma\vee(\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\to\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j})$ be a standard disjunction.
  1. (1)
  2. (2)
    if $\mathsf{S}\notin\mathsf{X}$, then $\mathbb{I}^{\prime}\neq\emptyset$;
  3. (3)
Example 2 (Neat sets of negative indices).
Assume $\mathsf{X}=\mathsf{D}$ and $\mathtt{AG}=\{a,b\}$. Consider the standard disjunction
$\bot\lor\big{(}\big{(}\langle\hskip-2.0pt[\mathrm{A}_{-1}]\hskip-2.0pt\rangle\phi_{-1}\land\langle\hskip-2.0pt[\mathrm{A}_{-2}]\hskip-2.0pt\rangle\phi_{-2}\land\langle\hskip-2.0pt[\mathrm{A}_{-3}]\hskip-2.0pt\rangle\phi_{-3}\big{)}\rightarrow\big{(}\langle\hskip-2.0pt[\mathrm{B}_{0}]\hskip-2.0pt\rangle\psi_{0}\lor\langle\hskip-2.0pt[\mathrm{B}_{1}]\hskip-2.0pt\rangle\psi_{1}\big{)}\big{)}$
with respect to $\gamma=\bot$, $\mathbb{I}=\{-1,-2,-3\}$ and $\mathbb{J}=\{0,1\}$, given in Example 1, where $\mathrm{A}_{-1}=\emptyset$, $\phi_{-1}=\top$, $\mathrm{A}_{-2}=\{b\}$, $\phi_{-2}=q$, $\mathrm{A}_{-3}=\{a\}$, $\phi_{-3}=p$, $\mathrm{B}_{0}=\mathtt{AG}$, $\psi_{0}=\bot$, $\mathrm{B}_{1}=\mathtt{AG}$, and $\psi_{1}=p\land q$.
It can be verified the following subsets of $\mathbb{I}$ are $\mathsf{D}$-neat: $\mathbb{I}_{1}=\{-1\},\mathbb{I}_{2}=\{-2\},\mathbb{I}_{3}=\{-3\},\mathbb{I}_{4}=\{-1,-2\}$, and $\mathbb{I}_{5}=\{-1,-3\}$.
Let $\gamma\vee(\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\to\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j})$ be a standard disjunction. Define $\mathbb{J}_{0}=\{j\in\mathbb{J}\mid\mathrm{B}_{j}=\mathtt{AG}\}$, called the set of basic positive indices.
@@ 3.5 Downward validity lemma
Let $\mathsf{X}\in\mathrm{ES}$ and $\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\to\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j})$ be a standard disjunction. Then, if $\models_{{\mathsf{MCL}+\mathsf{X}}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\to\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j})$, then the following condition, called the $\mathsf{X}$-validity-reduction condition, is met:
@@
The following example illustrates this lemma.
Example 3.
Assume $\mathsf{X}=\mathsf{D}$ and $\mathtt{AG}=\{a,b\}$. Consider the standard disjunction
$\bot\lor\big{(}\big{(}\langle\hskip-2.0pt[\mathrm{A}_{-1}]\hskip-2.0pt\rangle\phi_{-1}\land\langle\hskip-2.0pt[\mathrm{A}_{-2}]\hskip-2.0pt\rangle\phi_{-2}\land\langle\hskip-2.0pt[\mathrm{A}_{-3}]\hskip-2.0pt\rangle\phi_{-3}\big{)}\rightarrow\big{(}\langle\hskip-2.0pt[\mathrm{B}_{0}]\hskip-2.0pt\rangle\psi_{0}\lor\langle\hskip-2.0pt[\mathrm{B}_{1}]\hskip-2.0pt\rangle\psi_{1}\big{)}\big{)}$
with respect to $\gamma=\bot$, $\mathbb{I}=\{-1,-2,-3\}$ and $\mathbb{J}=\{0,1\}$, given in Example 1, where $\mathrm{A}_{-1}=\emptyset$, $\phi_{-1}=\top$, $\mathrm{A}_{-2}=\{b\}$, $\phi_{-2}=q$, $\mathrm{A}_{-3}=\{a\}$, $\phi_{-3}=p$, $\mathrm{B}_{0}=\mathtt{AG}$, $\psi_{0}=\bot$, $\mathrm{B}_{1}=\mathtt{AG}$, and $\psi_{1}=p\land q$.
As mentioned in Example 2, the following subsets of $\mathbb{I}$ are $\mathsf{D}$-neat: $\mathbb{I}_{1}=\{-1\},\mathbb{I}_{2}=\{-2\},\mathbb{I}_{3}=\{-3\},\mathbb{I}_{4}=\{-1,-2\}$, and $\mathbb{I}_{5}=\{-1,-3\}$. Note $\mathbb{J}_{0}=\{0,1\}$.
By this lemma, if $\models_{\mathsf{MCL}+\mathsf{D}}\phi$, then one of the following holds:
  1. (1)
    $\models_{\mathsf{MCL}+\mathsf{D}}\gamma$, that is, $\models_{\mathsf{MCL}+\mathsf{D}}\bot$;
  2. (2)
    $\bigcup_{i\in\mathbb{I}_{1}}\mathrm{A}_{i}\subseteq\mathrm{B}_{0}$ and $\models\bigwedge_{i\in\mathbb{I}_{1}}\phi_{i}\rightarrow(\psi_{0}\land\bigwedge_{j\in\mathbb{J}_{0}}\psi_{j})$, that is, $\emptyset\subseteq\mathtt{AG}$ and $\models_{\mathsf{MCL}+\mathsf{D}}\top\rightarrow(\bot\land\bot\land(p\land q))$;
  3. (3)
    $\bigcup_{i\in\mathbb{I}_{2}}\mathrm{A}_{i}\subseteq\mathrm{B}_{0}$ and $\models\bigwedge_{i\in\mathbb{I}_{2}}\phi_{i}\rightarrow(\psi_{0}\land\bigwedge_{j\in\mathbb{J}_{0}}\psi_{j})$, that is, $\{b\}\subseteq\mathtt{AG}$ and $\models_{\mathsf{MCL}+\mathsf{D}}q\rightarrow(\bot\land\bot\land(p\land q))$;
  4. (4)
    $\bigcup_{i\in\mathbb{I}_{3}}\mathrm{A}_{i}\subseteq\mathrm{B}_{0}$ and $\models\bigwedge_{i\in\mathbb{I}_{3}}\phi_{i}\rightarrow(\psi_{0}\land\bigwedge_{j\in\mathbb{J}_{0}}\psi_{j})$, that is, $\{a\}\subseteq\mathtt{AG}$ and $\models_{\mathsf{MCL}+\mathsf{D}}p\rightarrow(\bot\land\bot\land(p\land q))$;
  5. (5)
    $\bigcup_{i\in\mathbb{I}_{4}}\mathrm{A}_{i}\subseteq\mathrm{B}_{0}$ and $\models\bigwedge_{i\in\mathbb{I}_{4}}\phi_{i}\rightarrow(\psi_{0}\land\bigwedge_{j\in\mathbb{J}_{0}}\psi_{j})$, that is, $\{b\}\subseteq\mathtt{AG}$ and $\models_{\mathsf{MCL}+\mathsf{D}}(\top\land q)\rightarrow(\bot\land\bot\land(p\land q))$;
  6. (6)
    $\bigcup_{i\in\mathbb{I}_{5}}\mathrm{A}_{i}\subseteq\mathrm{B}_{0}$ and $\models\bigwedge_{i\in\mathbb{I}_{5}}\phi_{i}\rightarrow(\psi_{0}\land\bigwedge_{j\in\mathbb{J}_{0}}\psi_{j})$, that is, $\{a\}\subseteq\mathtt{AG}$ and $\models_{\mathsf{MCL}+\mathsf{D}}(\top\land p)\rightarrow(\bot\land\bot\land(p\land q))$;
  7. (7)
    $\bigcup_{i\in\mathbb{I}_{1}}\mathrm{A}_{i}\subseteq\mathrm{B}_{1}$ and $\models\bigwedge_{i\in\mathbb{I}_{1}}\phi_{i}\rightarrow(\psi_{1}\land\bigwedge_{j\in\mathbb{J}_{0}}\psi_{j})$, that is, $\emptyset\subseteq\mathtt{AG}$ and $\models_{\mathsf{MCL}+\mathsf{D}}\top\rightarrow((p\land q)\land\bot\land(p\land q))$;
  8. (8)
    $\bigcup_{i\in\mathbb{I}_{2}}\mathrm{A}_{i}\subseteq\mathrm{B}_{1}$ and $\models\bigwedge_{i\in\mathbb{I}_{2}}\phi_{i}\rightarrow(\psi_{1}\land\bigwedge_{j\in\mathbb{J}_{0}}\psi_{j})$, that is, $\{b\}\subseteq\mathtt{AG}$ and $\models_{\mathsf{MCL}+\mathsf{D}}q\rightarrow((p\land q)\land\bot\land(p\land q))$;
  9. (9)
    $\bigcup_{i\in\mathbb{I}_{3}}\mathrm{A}_{i}\subseteq\mathrm{B}_{1}$ and $\models\bigwedge_{i\in\mathbb{I}_{3}}\phi_{i}\rightarrow(\psi_{1}\land\bigwedge_{j\in\mathbb{J}_{0}}\psi_{j})$, that is, $\{a\}\subseteq\mathtt{AG}$ and $\models_{\mathsf{MCL}+\mathsf{D}}p\rightarrow((p\land q)\land\bot\land(p\land q))$;
  10. (10)
    $\bigcup_{i\in\mathbb{I}_{4}}\mathrm{A}_{i}\subseteq\mathrm{B}_{1}$ and $\models\bigwedge_{i\in\mathbb{I}_{4}}\phi_{i}\rightarrow(\psi_{1}\land\bigwedge_{j\in\mathbb{J}_{0}}\psi_{j})$, that is, $\{b\}\subseteq\mathtt{AG}$ and $\models_{\mathsf{MCL}+\mathsf{D}}(\top\land q)\rightarrow((p\land q)\land\bot\land(p\land q))$;
  11. (11)
    $\bigcup_{i\in\mathbb{I}_{5}}\mathrm{A}_{i}\subseteq\mathrm{B}_{1}$ and $\models\bigwedge_{i\in\mathbb{I}_{5}}\phi_{i}\rightarrow(\psi_{1}\land\bigwedge_{j\in\mathbb{J}_{0}}\psi_{j})$, that is, $\{a\}\subseteq\mathtt{AG}$ and $\models_{\mathsf{MCL}+\mathsf{D}}(\top\land p)\rightarrow((p\land q)\land\bot\land(p\land q))$.
We now show the downward validity lemma.
@@
Assumptions
Assume the $\mathsf{X}$-validity-reduction condition is not met. Then, it can be verified that the following condition, called the $\mathsf{X}$-satisfiability-reduction condition, is met:
@@
Let $\gamma^{\prime}$ be an $\mathsf{X}$-satisfiable elementary conjunction, equivalent to $\neg\gamma$. It suffices to show the standard conjunction $\gamma^{\prime}\wedge\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\wedge\bigwedge_{j\in\mathbb{J}}\neg\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$ is $\mathsf{X}$-satisfiable.
In the sequel, first, we define a blueprint, from which we can get a pointed $\mathsf{X}$-model; second, we show the pointed $\mathsf{X}$-model satisfies $\gamma^{\prime}\wedge\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\wedge\bigwedge_{j\in\mathbb{J}}\neg\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$.
To define the blueprint, we need two functions: $\mathtt{support}$ and $\mathtt{impeach}$.
Two functions $\mathtt{support}$ and $\mathtt{impeach}$
@@
Definition of a blueprint $\mathtt{BP}$
Define a blueprint $\mathtt{BP}=(\mathtt{AC}_{0},\mathtt{list}_{\mathtt{AG}})$ as follows:
@@
    $\begin{cases}\Big{\{}\bigwedge_{i\in\mathtt{support}({\sigma_{\mathtt{AG}}})}\phi_{i}\wedge\neg\psi_{j}\mid j\in\mathbb{J}\text{ and }\bigcup_{i\in\mathtt{support}({\sigma_{\mathtt{AG}}})}\mathrm{A}_{i}\subseteq\mathrm{B}_{j}\Big{\}}&\parbox[t]{110.00017pt}{ \text{if $\mathtt{support}({\sigma_{\mathtt{AG}}})$ is $\mathsf{X}$-neat} }\\ \emptyset&\text{otherwise}\end{cases}$
@@
[…]
    otherwise
[…]
The bluepint $\mathtt{BP}$ is $\mathsf{X}$-regular
@@
    • Assume $\mathsf{D}\notin\mathsf{X}$. Then there is $j\in\mathbb{J}$ such that $\chi=\bigwedge_{i\in\mathtt{support}({\sigma_{\mathtt{AG}}})}\phi_{i}\wedge\neg\psi_{j}$, where $\mathtt{support}({\sigma_{\mathtt{AG}}})$ is $\mathsf{X}$-neat and $\bigcup_{i\in\mathtt{support}({\sigma_{\mathtt{AG}}})}\mathrm{A}_{i}\subseteq\mathrm{B}_{j}$. By the $\mathsf{X}$-satisfiability-reduction condition, $\chi$ is $\mathsf{X}$-satisfiable.
@@
      • Assume $\mathtt{support}({\sigma_{\mathtt{AG}}})=\emptyset$. We want to show that this case is impossible. By Claim (3), $\mathbb{I}=\emptyset$. Therefore, $\mathtt{support}(\lambda_{\mathtt{AG}})=\emptyset$ for all $\lambda_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$. Note ${\sigma_{\mathrm{C}}}\in\mathtt{PJA}_{\mathrm{C}}$. Then $\mathtt{list}_{\mathrm{C}}(\sigma_{\mathrm{C}})\neq\emptyset$. Then there is $\lambda_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\lambda_{\mathtt{AG}}$ and $\mathtt{list}_{\mathtt{AG}}(\lambda_{\mathtt{AG}})\neq\emptyset$. By Claim (4), $\mathtt{support}(\lambda_{\mathtt{AG}})$ is $\mathsf{X}$-neat. Note $\mathsf{S}\notin\mathsf{X}$. Then $\mathtt{support}(\lambda_{\mathtt{AG}})\neq\emptyset$. We have a contradiction.
@@
The standard conjunction $\gamma^{\prime}\wedge\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\wedge\bigwedge_{j\in\mathbb{J}}\neg\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$ is $\mathsf{X}$-satisfiable
By Theorem 1, there is a pointed $\mathsf{X}$-model $(\mathtt{M},s)$, where $\mathtt{M}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$, realizing $\mathtt{BP}$ and $\gamma^{\prime}$. We claim $\mathtt{M},s\Vdash\gamma^{\prime}\land\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\wedge\bigwedge_{j\in\mathbb{J}}\neg\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$.
We show this in two steps.
Step one
Let $i\in\mathbb{I}$. We want to show $\mathtt{M},s\Vdash\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}$. Let ${\sigma_{\mathrm{A}_{i}}}\in\mathtt{JA}_{\mathrm{A}_{i}}$ such that ${{\sigma_{\mathrm{A}_{i}}}(a)}=i$ for all $a\in\mathrm{A}_{i}$. It suffices to show $\sigma_{\mathrm{A}_{i}}\in\mathtt{PJA}_{\mathrm{A}_{i}}$ and ${\sigma_{\mathrm{A}_{i}}}\vartriangleright_{(\mathtt{M},s)}\phi_{i}$.
@@
Step two
Let $j\in\mathbb{J}$. We want to show $\mathtt{M},s\Vdash\neg\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$. Let ${\sigma_{\mathrm{B}_{j}}}\in\mathtt{PJA}_{\mathrm{B}_{j}}$. It suffices to show ${\sigma_{\mathrm{B}_{j}}}\ltimes_{(\mathtt{M},s)}\neg\psi_{j}$.
Note $\mathtt{list}_{\mathrm{B}_{j}}({\sigma_{\mathrm{B}_{j}}})\neq\emptyset$. Then there is $\lambda_{\mathtt{AG}}\in\mathtt{PJA}_{\mathtt{AG}}$ such that ${\sigma_{\mathrm{B}_{j}}}\subseteq\lambda_{\mathtt{AG}}$ and $\mathtt{list}_{\mathtt{AG}}(\lambda_{\mathtt{AG}})\neq\emptyset$. By Claim (4), $\mathtt{support}(\lambda_{\mathtt{AG}})$ is $\mathsf{X}$-neat. By Claim (2), $\mathtt{support}({\sigma_{\mathrm{B}_{j}}})\subseteq\mathtt{support}(\lambda_{\mathtt{AG}})$. We claim $\mathtt{support}({\sigma_{\mathrm{B}_{j}}})$ is $\mathsf{X}$-neat. It is easy to check that the first and third conditions given in Definition 16, the definition of neatness, hold for $\mathtt{support}({\sigma_{\mathrm{B}_{j}}})$.
Assume $\mathsf{S}\in\mathsf{X}$. Clearly, $\mathtt{support}({\sigma_{\mathrm{B}_{j}}})$ meets the second condition given in Definition 16. Then $\mathtt{support}({\sigma_{\mathrm{B}_{j}}})$ is $\mathsf{X}$-neat.
Assume $\mathsf{S}\notin\mathsf{X}$. Note $\mathtt{support}(\lambda_{\mathtt{AG}})$ is $\mathsf{X}$-neat. Then $\mathtt{support}(\lambda_{\mathtt{AG}})\neq\emptyset$. Then $\mathbb{I}\neq\emptyset$ and $1\in\mathbb{I}$. Note $\mathrm{A}_{-1}=\emptyset$. It is easy to check $1\in\mathtt{support}({\sigma_{\mathrm{B}_{j}}})$. Then the second condition given in Definition 16 holds for $\mathtt{support}({\sigma_{\mathrm{B}_{j}}})$. Then $\mathtt{support}({\sigma_{\mathrm{B}_{j}}})$ is $\mathsf{X}$-neat.
We consider two possible cases, respectively.
@@
  • Assume $\mathrm{A}_{i}\nsubseteq\mathrm{B}_{j}$. We want to show that this case is impossible. Let $\overline{b}\in\mathrm{A}_{i}-\mathrm{B}_{j}$. Then $\overline{b}\in\mathtt{AG}-\mathrm{B}_{j}$. Then ${{\sigma_{\mathtt{AG}}}(\overline{b})}=j$. Note $\overline{b}\in\mathrm{A}_{i}$. Then ${{\sigma_{\mathtt{AG}}}(\overline{b})}=i$. Then $j=i$. Note $i\in\mathbb{I}$ and $j\in\mathbb{J}$. Then $\mathbb{I}\cap\mathbb{J}\neq\emptyset$. We have a contradiction.